OpenStudy (anonymous):

find the roots of the equation x^2-4x=21 by factoring

OpenStudy (anonymous):

\[x^2-4x-21=0\] \[(x-7)(x+3)=0\] \[x-7=0\] \[x=7\] \[x+3=0\] \[x=-3\]